OpenStudy (anonymous):

Help with algebra 2. Question in comments, Plz help :3

OpenStudy (anonymous):

\[3^{x}=2\]

OpenStudy (anonymous):

i think it's \[\frac{ 2\log }{ 3\log }\]

OpenStudy (anonymous):

i meant log2/log3 not 2log/3log...